# Quantum geometric formulation of Brans-Dicke theory for Bianchi I spacetime

### v2

@@ V.2 The $\bar{\mu}_{B}$ quantization scheme
where $I$, $J$ and $K$ are non repeated indexes.
More explicilty, we begin by rewriting Eq. (45) as
@@
At this point, we apply the aforementioned alternative scheme, given by Eq. (54), in Eq. (V.2), which can now be suitably written such as to define an effective energy density $\rho_{\rm eff}$ as
$\frac{\rho_{\rm eff}}{\rho_{c}}$ $=$ $\frac{\gamma^{2}\Delta}{3}\left(\frac{\beta\dot{\phi}^{2}}{4}+\rho\phi\right)$ (58)
@@
Therefore, the effective energy density $\rho_{\rm eff}$, already defined by Eq. (52), is bounded from above by its maximal value $\rho_{c}$, given by $\rho_{c}=3/(\gamma^{2}\Delta)$. These results are in agreement with Ref. [53] in the isotropic limit.
Now, a maximal value of energy density implies a minimum value of spatial volume. Therefore, initializing the universe with a contracting phase, a big bounce is predicted in this scenario. This confirms the removal of an initial singularity in the framework of LQG owing to the fact that geometry is quantized and there exists a non-zero minimal area. Starting with an initially contracting phase, in a classical setup, the collapse is inevitable. Thus, once again, through this analysis we conclude that quantum geometric effects at the Planck scale provide repulsive forces and replace the big bang singularity by a non-singular quantum bounce even beyond GR, with the non-trivial spacetime, the Bianchi I universe.
For completeness, let us now write a closed form of the Friedmann equation and that can be derived using the current scheme. From Eqs. (LABEL:HubbleParameter) and (55), we obtain that
$\left(H+\frac{\dot{\phi}}{2\phi}\right)^{2}$ $=$ $\left\{\frac{1}{6\gamma\sqrt{\Delta}\phi}\left[\sin(\tilde{c}_{1}\bar{\mu}_{1})\cos(\tilde{c}_{2}\bar{\mu}_{2})+\sin(\tilde{c}_{2}\bar{\mu}_{2})\cos(\tilde{c}_{1}\bar{\mu}_{1})+\sin(\tilde{c}_{1}\bar{\mu}_{1})\cos(\tilde{c}_{3}\bar{\mu}_{3})+\sin(\tilde{c}_{3}\bar{\mu}_{3})\cos(\tilde{c}_{1}\bar{\mu}_{1})+\right.\right.$ (59)
@@
Inspired by Eq. (58), we define the directional effective energy density in the I-direction as
@@
$\left(H+\frac{\dot{\phi}}{2\phi}\right)^{2}$ $=$ $\left\{\frac{1}{6\phi}\left[\sqrt{\frac{1}{3}\rho_{\rm eff}^{(1)}\left(1-\frac{\rho_{\rm eff}^{(2)}}{\rho_{c}}\right)}+\sqrt{\frac{1}{3}\rho_{\rm eff}^{(2)}\left(1-\frac{\rho_{\rm eff}^{(1)}}{\rho_{c}}\right)}+\sqrt{\frac{1}{3}\rho_{\rm eff}^{(1)}\left(1-\frac{\rho_{\rm eff}^{(3)}}{\rho_{c}}\right)}+\right.\right.$ (61)
@@
VI Dynamics of the Models
In this section we present the results for the effective dynamics of the models considered in this work. We show the evolution of the scale factor and Hubble parameter for Bianchi LQC and BD Bianchi LQC models in each quantization scheme that has been discussed in the previous section.
In Figs. 1 and 2 we show the results for the directional scale factor $a_{I}$, $I=1,2,3$, in a Bianchi-I LQC model for the $\bar{\mu}_{A}$ and $\bar{\mu}_{B}$ schemes, given by Eqs. (41) and (54), respectively, with the parameters described in the figure captions. In Figs. 1(a) and 2(a), we show the results for the pure Bianchi LQC, while in Figs. 1(b) and 2(b) we have the corresponding results in the context of BDT. The initial conditions considered, which are described in the figure caption, are set at the contracting pre-bounce phase and at an instant we denote by $t=0$ (as will be the case for all the results presented in this Section). We can see from the figure that the behavior of the scale factor differs in each direction, as expected for Bianchi spacetimes. Especially, the behavior of $a_{1}$ is very different from $a_{2}$ and $a_{3}$ in both models considered. This type of dynamics is consistent with a Kasner-like phase, i.e., expansion followed by contraction in one direction and contraction followed by expansion in the others. It is important to observe that even though there appears contraction in one of the directions, we have checked explicitly that it never reaches zero (e.g., reaches a singularity) in that direction, but it reaches a nonnull value at very large times and then increases. The results shown in Figs. 1 and 2 are in agreement with the overall behavior obtained in Refs. [36] and [54] for the standard (non-BD, i.e. Einstein-Hilbert) Bianchi LQC case. From the numerical studies that we have performed, we have also verified that the inclusion of a massless scalar field does not change the results of the vacuum case significantly, although the exact behavior can depend on the initial conditions and on the choice of matter.
@@ VI Dynamics of the Models
Figure 1: The directional scale factors $a_{I}$, $I=1,2,3$, in a Bianchi-I LQC model (panel a) and for the BD Bianchi-I LQC model (panel b), for the initial conditions $c_{1}=0.6$, $c_{2}=-3.0$, $c_{3}=5.5$, $p_{1}=9$, $p_{2}=9.0001$, $p_{3}=8.9999$, and for the BD Bianchi-I LQC model, with the BD initial conditions $\phi=1$, $p_{\phi}=0$ and for $\omega=0.1$ for the BD parameter. Both plots correspond to the $\bar{\mu}_{A}$ scheme.
Figure 2: The directional scale factors $a_{I}$, $I=1,2,3$, in a Bianchi-I LQC model (panel a) and for the BD Bianchi-I LQC model (panel b), for the initial conditions $c_{1}=-10$, $c_{2}=-15$, $c_{3}=20$, $p_{1}=9$, $p_{2}=9.00001$, $p_{3}=8.99999$, and for the BD Bianchi-I LQC model, with the BD initial conditions $\phi=1$, $p_{\phi}=0$ and for $\omega=0.1$ for the BD parameter. Both plots correspond to the $\bar{\mu}_{B}$ scheme.
Besides of checking the behavior of the scale factor in each direction, the important quantity to investigate, in order to analyze the effective cosmological dynamics in these models, is the average scale factor and the average Hubble parameter (where the average is taken with respect to the three directions).
We show in Fig. 3 the behavior of the average scale factor for the Bianchi LQC (+BD) in the two quantization schemes. We can see that the evolution is different in each scheme, as previously predicted in the work of Ref. [54]. By comparing the results within the same quantization scheme, we can see that the slopes of the curves are different depending on whether the model is in the context of the Brans-Dicke theory or not. In order to further investigate the dynamics in each case, it is worth to investigate the evolution of the Hubble parameter.
In Fig. 4, we present the evolution of the Hubble parameter in each model considered. For the models in the context of the $\bar{\mu}_{A}$-scheme, we can observe the presence of a single bounce. On the other hand, the curves representing the models in the $\bar{\mu}_{B}$-scheme crosses the value $H=0$ more than once. By comparing each curve within the same quantization scheme, it becomes clear how the BD parameters affect the effective dynamics of the models. It is also possible to verify that the green (dashed) curve in Fig. 4, representing the dynamics of BD Bianchi LQC in the $\bar{\mu}_{A}$-scheme, shows an overall behavior consistent with the results obtained in Ref. [39].
Figure 3: The mean scale factor $a$. The initial conditions for Bianchi-I LQC model are $c_{1}=0.6$, $c_{2}=-3.0$, $c_{3}=5.5$, $p_{1}=9$, $p_{2}=9.0001$, $p_{3}=8.9999$, whereas for the BD Bianchi-I LQC model we used the same ones with the addition of $\phi=1$ and $p_{\phi}=0$. We chose the value $\omega=0.1$ for the BD parameter.
Figure 4: The mean Hubble parameters $H$. The initial conditions for Bianchi-I LQC model are $c_{1}=0.6$, $c_{2}=-3.0$, $c_{3}=5.5$, $p_{1}=9$, $p_{2}=9.0001$, $p_{3}=8.9999$, whereas for the BD Bianchi-I LQC model we used the same ones with the addition of $\phi=1$ and $p_{\phi}=0$. We chose the value $\omega=0.1$ for the BD parameter.
We conclude that, for the representative parameters and initial conditions considered, all the results obtained consistently show that a bounce occurs around $t\sim 0.1/M_{\rm pl}$, although the precise instant varies little within each model. The results present a behavior consistent with the ones previously obtained in the literature (for Bianchi LQC or BD LQC) in the respective limits. Although the detailed dynamics varies with the quantization scheme and the choice of initial conditions, the avoidance of the singularity is verified for all the cases considered.
@@
In this section we did not include the presence of stiff matter since we verified that it does not change the general results presented here, as the qualitative behavior shown in the figures remains the same. We have also explicitly verified that the values of the parameters and initial conditions chosen are quite generic in the sense that, by changing them to other reasonable values, we obtain the same generic behavior, such that main conclusions and results discussed above are maintained.
VII Conclusions
@@
In this paper we have derived the quantum-corrected effective background dynamics for BDT in the Bianchi I model in the Jordan’s frame in the framework of LQG. We have cast the theory in terms of the Yang-Mills phase space variables suitable for loop quantization and we have presented the quantum-corrected effective dynamics that resolve the spacetime singularity. The quantization is made by considering two different schemes, as given by Eqs. (41) and (54), respectively. The former set provides lower bounds on the scale factor, Eq. (53), proving the existence of a non-singular quantum behavior of spacetime. On the other hand, the latter scheme gives us an upper bound to the effective energy density, i.e., the critical density $\rho_{c}$, implying a minimal value of spatial volume. Additionally, we have also derived the quantum corrected Friedmann equation for this model inspired by the idea of a directional energy density. We have explicitly shown that the effective dynamics of the Bianchi I universe must undergo a smooth transition from a contracting phase to a expanding phase through the quantum bounce. Our analytical results, therefore, show robustness of the singularity resolution and the consistency of the effective dynamics beyond general relativity.
@@
As a perspective of future study in the context of the model here studied, it would be of interest to explore any possible observational features it might lead and that can be detected in the forthcoming CMB experiments. An extension of the analysis like the one performed in earlier studies in the context of LQC for homogeneous and isotropic models [55, 56, 57, 58, 59] would be of interest. Likewise, the analysis of the perturbations in the model presented here and the possibility of an inflationary phase following the quantum bounce could be eventually done. Concerning any possible observational signatures in LQC scenarios with an anisotropic background, one can mention, for instance, the work of Ref. [28], where an analysis of the possible observational features in such models was performed. Hence, an extension of such analysis of perturbations in the context of the BDT anisotropic LQC will be necessary in order to elaborate the observable predictions of the model. Issues related to particle production due to the bounce can also be important [60, 61, 62, 63, 64, 65, 66, 67, 68] and their analysis and consequences in the context of anisotropic loop quantum BDT would be worthwhile to be pursued in the future as well.

### v3

@@ IV.2 The $\bar{\mu}_{B}$ quantization scheme
where $I$, $J,$ and $K$ are nonrepeated indexes. The corresponding effective Hamiltonian takes the same form as that given by Eq. (38). Following the same steps provided in the last subsection, we can obtain the corresponding Hamiltonian equations. In particular, it can be shown that
@@
At this point, we apply the aforementioned alternative scheme, given by Eq. (53), in Eq. (IV.2), and define an effective energy density $\rho_{\rm eff}$ as
$\frac{\rho_{\rm eff}}{\rho_{c}}$$\equiv$ $\frac{\gamma^{2}\Delta}{3}\left(\frac{\beta\dot{\phi}^{2}}{4}+\rho\phi\right)$ (57)
@@
Clearly, such defined effective energy density is bounded from above by its maximal value $\rho_{c}$, given by $\rho_{c}=3/(\gamma^{2}\Delta)$. These results are in agreement with Ref. [74] in the isotropic limit.
Now, a maximal value of energy density implies a minimum value of spatial volume. Therefore, initializing the Universe with a contracting phase, a quantum bounce is predicted in this scenario. This confirms the removal of an initial singularity in the framework of LQG owing to the fact that geometry is quantized and there exists a nonzero minimal area gap. Starting with an initially contracting phase, in a classical setup, the collapse will inevitably lead to a spacetime singularity. Thus, once again, through this analysis we conclude that quantum geometric effects at the Planck scale provide repulsive forces and replace classical spacetime singularities by nonsingular quantum bounces.
For completeness, let us now write a closed form of the modified Friedmann equation and that can be derived using the current scheme. From Eqs. (LABEL:HubbleParameter) and (54), we obtain that
$\left(H+\frac{\dot{\phi}}{2\phi}\right)^{2}$ $=$ $\left\{\frac{1}{6\gamma\sqrt{\Delta}\phi}\left[\sin(\tilde{c}_{1}\bar{\mu}_{1})\cos(\tilde{c}_{2}\bar{\mu}_{2})+\sin(\tilde{c}_{2}\bar{\mu}_{2})\cos(\tilde{c}_{1}\bar{\mu}_{1})+\sin(\tilde{c}_{1}\bar{\mu}_{1})\cos(\tilde{c}_{3}\bar{\mu}_{3})+\sin(\tilde{c}_{3}\bar{\mu}_{3})\cos(\tilde{c}_{1}\bar{\mu}_{1})\right.\right.$ (58)
@@
Inspired by Eq. (57), we define the directional effective energy density in the I direction as
@@
$\left(H+\frac{\dot{\phi}}{2\phi}\right)^{2}$ $=$ $\left\{\frac{1}{6\phi}\left[\sqrt{\frac{1}{3}\rho_{\rm eff}^{(1)}\left(1-\frac{\rho_{\rm eff}^{(2)}}{\rho_{c}}\right)}+\sqrt{\frac{1}{3}\rho_{\rm eff}^{(2)}\left(1-\frac{\rho_{\rm eff}^{(1)}}{\rho_{c}}\right)}+\sqrt{\frac{1}{3}\rho_{\rm eff}^{(1)}\left(1-\frac{\rho_{\rm eff}^{(3)}}{\rho_{c}}\right)}\right.\right.$ (60)
@@
V Dynamics of Bianchi I Universe
In this section we present our numerical results for the effective dynamics of the models considered in the last section. We show the evolution of the scale factors and Hubble parameters of the Bianchi I Universe for LQC and BDT in each of the quantization schemes. To be able to compare our results with the ones obtained in Ref. [47] in the framework of LQC, here we choose the same initial conditions as those used by the authors of that reference. On the other hand, in the framework of BDT, we choose the initial condition for the triad and the connection variables as close as possible to those in LQC. In particular, in all of our numerical results for LQC, we choose the initial conditions $c_{1}=-0.13$, $c_{2}=-0.12$, $p_{1}=10^{3}$, $p_{2}=2\times 10^{3}$, $p_{3}=3\times 10^{3}$, while $c_{3}$ is determined from the Hamiltonian constraint, Eq. (55), with $(\phi,\dot{\phi},\rho)=(1,0,0)$.
With the above in mind, in Figs. 1 and 2 we show our numerical results for the directional scale factors $a_{I}\;(I=1,2,3)$ of the Bianchi-I spacetimes, in both LQC and BDT, for the $\bar{\mu}_{A}$ and $\bar{\mu}_{B}$ schemes, given, respectively by Eqs. (37) and (53), with the parameters described in the figure captions. Each figure has four panels for each scheme. Figures 1(a) and 2(a) shows the results for the LQC model, while Figs. 1(b)-1(d) and 2(b)-2(d) show the results for the BDT models. Notice that we use initial conditions for $\dot{\phi}$ instead of $\pi_{\phi}$ because it is more transparent when considering the limit $\mathrm{BDT}\to\mathrm{LQC}$, which happens for $(\phi,\dot{\phi})\to(1,0)$.
@@ V Dynamics of Bianchi I Universe
Figure 1: The evolution of the directional scale factors $a_{I}\;(I=1,2,3)$ of the Bianchi-I spacetimes, where (a) is for LQC and (b) - (d) are for BDT. The initial conditions for LQC are $c_{1}=-0.13$, $c_{2}=-0.12$, $p_{1}=10^{3}$, $p_{2}=2\times 10^{3}$, $p_{3}=3\times 10^{3}$. For BDT, we adopt the same initial conditions as those of LQC with the addition of $(\phi,\dot{\phi})=(1.5,0)$ for (b), $(\phi,\dot{\phi})=(1.0,5.5\times 10^{-5})$ for Panel (c), and $(\phi,\dot{\phi})=(1.0,-5.5\times 10^{-5})$ for (d), respectively, and set $\omega=2\times 10^{5}$ [58] for the BD parameter in all these three last. All plots correspond to the $\bar{\mu}_{A}$ scheme.
Figure 2: The same as in Fig. 1, but for the $\bar{\mu}_{B}$ scheme.
In Figs. 1(a) and 2(a) we show the results for the directional scale factors $a_{I}\;(I=1,2,3)$ of the Bianchi-I spacetime in the framework of LQC for the $\bar{\mu}_{A}$ and $\bar{\mu}_{B}$ schemes, respectively. Comparing them with the corresponding ones given in Ref. [47], one can see clearly the consistency among these figures within the numerical errors. As a result, the numerical code used in the current paper and the one used in Ref. [47] are all reliable and give the same results.
Then, comparing Fig. 1(a) with Fig. 1(b), we can see that the deviations of BDT from that of LQC are simply the delay of the bounce in each direction, while the main qualitative behavior is preserved. This is because in the present case the initial condition of $\dot{\phi}$ is chosen to be zero, as that in LQC, so at the beginning the BDT effects are negligible. This type of dynamics is consistent with a Kasner-like phase, i.e., expansion followed by contraction in one direction and contraction followed by expansion in the others. It is important to note that even though there appears contraction in one of the three directions, we have checked explicitly that it never reaches zero (e.g., reaches a spacetime singularity) in that direction, but it always reaches asymptotically a nonzero value. Similar behavior has been also observed previously in LQC, see, for example, Refs. [29] [30]  [47]. From the numerical studies that we have performed, we have also verified that the inclusion of a massless scalar field $\Phi$ does not change the results significantly, as long as $\dot{\phi}(t_{i})=0$, where $t_{i}$ denotes the initial time.
However, from Fig. 1(c) we can see that once the initial condition of $\dot{\phi}$ is different from zero, the deviations become significant. In particular, all the three directional scale factors are increasing after the bounce (at $M_{\rm Pl}t_{B}\simeq 33$) and soon after the bounce they become comparable with their initial values. This is sharply in contrast to the LQC case (first seen in Ref. [47]), in which one of the three directional scale factors is always decreasing, and only two of them are increasing after the bounce. Here we see that this is also true for both $\mu_{A}$ and $\mu_{B}$ schemes when $\dot{\phi}=0$, as one can see from Figs. 1(b) and 2(b). On the other hand, the difference between Figs. 1(c) and 1(d) is the choice of the sign of $\dot{\phi}$ at the initial moment. In fact, in Fig. 1(c) we chose $(\phi,\dot{\phi})=(1.0,5.5\times 10^{-5})$, while in Fig. 1(d) we chose $(\phi,\dot{\phi})=(1.0,-5.5\times 10^{-5})$. The results for both cases show that the directional scale factors remain the same. This can be seen from Eq. (49), in which its right-hand side is a constant, while its left-hand side is independent of the signs of $\dot{\phi}$, as one can see from Eq. (50). Recall that $p_{I}\equiv l_{0}^{2}\,a_{J}a_{K}$ for nonrepeated indexes $I,\,J,\,K$. Similar results as the ones for the $\mu_{A}$ scheme are also seen for the $\mu_{B}$ schemes. Comparing Fig. 2(a) with Fig. 2(b), we can see that the directional scale factors behave quite similar, modulated the simple delay of the bounce in each direction. In particular, only two of the three directional scale factors are increasing after the bounce, while the third one is always decreasing, which is again consistent as the results observed in Ref. [47]. Again, this is because of the choice of the initial condition $\dot{\phi}(t_{i})=0$, which is initially identical to those in the LQC case. However, once $\dot{\phi}(t_{i})\neq 0$, as shown explicitly in Figs. 2(c) and 2(d), the deviations become large. Remarkably, all the three directional scale factors are increasing right after the bounce and soon become comparable with their initial values. That is, if the evolution of the spacetime starts from a classical point, it will collapse to a minimal but nonzero volume, where the quantum bounce occurs and replaces the classical spacetime singularity by a finite quantum bounce. Afterwards, the spacetime starts to expand in each of the three directions, and soon approaches to a size that is comparable with the initial one. This is true not only in terms of volumes and areas but also in terms of the size in each of the three directions. This is dramatically different from that observed in LQC for the $\mu_{A}$ scheme [47].
To understand the properties of the models further, in Figs. 3 and  4, we show the numerical results for the average scale factor $a$ and Hubble parameter $H$, defined by Eq. (45), for both of the $\mu_{A}$ and $\mu_{B}$ schemes. In each of these figures, the initial conditions are chosen the same as the corresponding ones given in Figs. 1 and 2, respectively, for the $\mu_{A}$ and $\mu_{B}$ schemes. In particular, Figs.  3 and 4 show similar behaviors for the LQC models as presented in Refs. [30, 47], while for the BDT models, the effects of the BD scalar $\phi$ depend on the choice of the initial conditions. When $\dot{\phi}(t_{i})=0$, the effects with respect to LQC are simply to delay the bounce, and make the increase of the scale factor slower, while for the initial conditions with $\dot{\phi}(t_{i})\neq 0$, the effects are opposite, that is, the bounce happens earlier, and the average scale factor increases faster after the bounce. Again, such behaviors do not depend on the signs of $\dot{\phi}(t_{i})$ for the same reasons as explained above for the directional scale factors $a_{I}$’s. Hence, the results for the same absolute value of $\dot{\phi}(t_{i})$ are overlapping.
Note that an overall behavior of the average Hubble parameter is also consistent with the results obtained earlier in Ref. [54], although the effects of the initial conditions on the shapes of $a$ and $H$ were not considered there.
In addition, in all the figures we have plotted the curves only up to $M_{\rm Pl}t\lesssim 300$, but we have found that the asymptotical behavior of the curves remains the same. In fact, we have run our codes up to $M_{\rm Pl}t\simeq 10^{6}$ and found no changes. So, their asymptotical behavior remains the same.
Figure 3: The average scale factor $a$ of the Bianchi I spacetime for the $\bar{\mu}_{A}$ [(a)] and $\bar{\mu}_{B}$ [(b)] schemes, where the LQC model is presented by the black curve and the BDT models are presented by the blue, yellow and red curves, respectively for the choices, $(\phi,\dot{\phi})=(1.5,0)$, $(\phi,\dot{\phi})=(1.0,5.5\times 10^{-5})$, and $(\phi,\dot{\phi})=(1.0,-5.5\times 10^{-5})$. The other initial conditions are the same as those given in Fig. 1 for the $\mu_{A}$ scheme, and those given in Fig. 2 for the $\mu_{B}$ scheme.
Figure 4: The average Hubble parameters $H$ of the Bianchi I spacetime for the $\bar{\mu}_{A}$ scheme [(a)] and the $\bar{\mu}_{B}$ scheme [(b)], where the LQC model is presented by the black curve and the BDT models are presented by the blue, yellow and red curves, respectively, for the choices, $(\phi,\dot{\phi})=(1.5,0)$, $(\phi,\dot{\phi})=(1.0,5.5\times 10^{-5})$, and $(\phi,\dot{\phi})=(1.0,-5.5\times 10^{-5})$. The other initial conditions are the same as those given in Fig. 1 for the $\mu_{A}$ scheme and Fig. 2 for the $\mu_{B}$ scheme.
Analysis of these figures leads us to conclude that, for the representative parameters and initial conditions considered, all results consistently demonstrate that the classical singularity is replaced by a regular quantum bounce. In the absence of BD effects, this bounce occurs around $t=65/M_{\rm Pl}$. However, the BD effects can shift the bounce to earlier or later times depending on the initial conditions of the BD scalar $\phi(t_{i})$ and its derivative $\phi(t_{i})$, where $t_{i}$ denotes the initial time. Furthermore, the BD effects prevent the suppression of the directional scale factors after the bounce, a stark contrast to what was previously seen in the LQC case [30, 47].
@@
Finally, we note that we have explicitly verified that the values of the parameters and initial conditions chosen are quite generic in the sense that, by changing them to other reasonable values, we obtain similar behaviors, so our main conclusions and results are robust with different initial conditions. Also, in this section we do not include plots for the cases in which the massless scalar field $\Phi$ (which is equivalent to a stiff fluid) is present, as we find that it does not change the general results presented here, and the qualitative behavior also remains the same.
VI Conclusions
@@
In this paper we have derived the quantum-corrected effective background dynamics for BDT in the Bianchi I model in the Jordan’s frame in the framework of LQG. We have cast the theory in terms of the Yang-Mills phase space variables suitable for loop quantization and we have presented the quantum-corrected effective dynamics that resolve the spacetime singularity. The quantization is made by considering two different schemes, as given by Eqs. (37) and (53), respectively. The former set provides lower bounds on the scale factor, Eq. (52), proving the existence of a nonsingular quantum behavior of spacetime. On the other hand, the latter scheme gives us an upper bound to the effective energy density, i.e., the critical density $\rho_{c}$, implying a minimal value of spatial volume. Additionally, we have also derived the quantum corrected Friedmann equation for this model inspired by the idea of a directional energy density. We have explicitly shown that the effective dynamics of the Bianchi I Universe must undergo a smooth transition from a contracting phase to a expanding phase through the quantum bounce. Our analytical results, therefore, show robustness of the singularity resolution and the consistency of the effective dynamics beyond general relativity.
In addition, we have also found different results obtained in BDT from those obtained in LQC. In particular, we have found that all the three directional scale factors are generically increasing after the quantum bounce and soon reach to their initial values, as can be seen from Figs. 1(c) and 1(d) for the $\mu_{A}$ scheme and in Figs. 2(c) and 2(d) for the $\mu_{B}$ scheme. This is sharply in contrast to the LQC case, in which one of the three directional scale factors is always decreasing after the bounce, and only two of them are increasing, as can be seen from Figs. 1(a) and 2(a) in both schemes.
As a perspective of future study in the context of the model considered here, it would be of interest to explore any possible observational features it might lead and that can be detected in the forthcoming cosmic microwave background experiments. An extension of the analysis, like the one performed in earlier studies in the context of LQC for homogeneous and isotropic models [75, 76, 77, 78, 79], would be of great interest.
@@
Likewise, the analysis of the perturbations in the model presented here and the possibility of an inflationary phase following the quantum bounce could be eventually done. Concerning any possible observational signatures in LQC scenarios with an anisotropic background, one can mention, for instance, the work of Ref. [51], where an analysis of the possible observational features in such models was performed. Hence, an extension of such analysis of perturbations in the context of the BDT anisotropic LQC will be necessary in order to elaborate the observable predictions of the model.
Issues related to particle production due to the bounce can also be important [80, 81, 82, 83, 84, 85, 86, 87, 88] and their analysis and consequences in the context of anisotropic loop quantum BDT would be worthwhile to be pursued in the future as well.